You have just made your first $5,500 contribution to your retirement account. Assume you earn a return of 10 percent per year and make no additional contributions. a. What will your account be worth when you retire in 45 years? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) b. What if you wait 10 years before contributing? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

a.

A=$5500*(1.1)^45

=$5500*72.89048369

=$400,897.66(Approx).

b.

A=$5500*(1.1)^35

=$5500*28.10243685

which is equal to

=$154,563.40(Approx).
Answer #2

SOLUTION :


a.  $400897.66 (ANSWER)


Work :


FV available on retirement after 44 years :

= Amount invested ( 1 + r)^n 

= 5500 (1 + 0.10)^45

= 400897.66 ($)  


b.  $154,563.40 (ANSWER)


Work :


If deposit of $5,500 is made after 10 years from now,

FV available at retirement after 44 years :

= Amount invested ( 1 + r)^n 

= 5500 (1 + 0.10)^(45 - 10)

= 154563.40 ($)  

answered by: Tulsiram Garg